A particle is being rotated in a centrifuge which has a radius of 5.0 m. If the 

particle’s centripetal acceleration is 4 g, determine its speed. What is the time period 

of its motion?


1
Expert's answer
2021-01-19T07:11:48-0500

a) By the definition of the centripetal acceleration, we have:


"a_c=\dfrac{v^2}{r},""v=\sqrt{a_cr}=\sqrt{4\cdot9.8\ \dfrac{m}{s^2}\cdot5.0\ m}=14\ \dfrac{m}{s}."

b) Let's first find the angular velocity from the formula:


"v=\omega r,""\omega=\dfrac{v}{r}=\dfrac{14\ \dfrac{m}{s}}{5.0\ m}=2.8\ \dfrac{rad}{s}."

Finally, we can find the time period of particle's motion from the formula:


"T=\dfrac{2\pi}{\omega}=\dfrac{2\pi}{2.8\ \dfrac{rad}{s}}=2.24\ s."

Answer:

a) "v=14\ \dfrac{m}{s}."

b) "T=2.24\ s."